OpenStudy (asiajenell):

MEDAL Solving Exponential equations - Algebra 2 6^2*(1/216)^-3n=1/216

OpenStudy (chiunlee12):

Did you learn log or ln yet?

OpenStudy (asiajenell):

yes

OpenStudy (chiunlee12):

Ah okay

OpenStudy (chiunlee12):

First, you would simplify the problem even further so: 36 * (1/216)^(-3n) = 1/216

OpenStudy (chiunlee12):

ok from there, you divide both sides by 36: (1/216)^(-3n) = 1/7776

OpenStudy (chiunlee12):

and I think here, you can change all the numbers to the base of 6

OpenStudy (chiunlee12):

So, \[(\frac{ 1 }{ 6^{3} })^{-3n}= \frac{ 1 }{ 6 }^{5}\]

OpenStudy (chiunlee12):

so according to exponential rules that would give: (1/6)^(-3n+3)=(1/6)^5

OpenStudy (asiajenell):

i was suppose to solve for n

OpenStudy (chiunlee12):

In this case, I don't even think you need to use log because you can just say

OpenStudy (chiunlee12):

Yeah, Im getting there, sorry if I'm too slow

OpenStudy (chiunlee12):

so

OpenStudy (asiajenell):

its okay

OpenStudy (chiunlee12):

You simplify to -3n+3=5

OpenStudy (chiunlee12):

And if you solve for n there, you can get the answer :)

OpenStudy (asiajenell):

thank you!!

OpenStudy (chiunlee12):

No problem